# Combined Pre-Supernova Alert System with KamLAND and Super-Kamiokande

### v2

@@ 5.6 Discussion
The results shown above suggest that the two detectors, KamLAND and SK, have different advantages in per-SN neutrino detection. For KamLAND, the low background rate is an advantage in terms of resolving a small signal. As shown in Figure 6 (a), for a Betelgeuse-like pre-SN candidate, the expected number of signal events exceed the background count even tens of hours prior to the CCSN. As a result, KamLAND can provide a warning for nearby supernovae earlier than SK. On the other hand, the number of signal events in KamLAND is limited by the target mass, making it hard to reach stars at far distances. SK is sensitive to CCSN candidates further away from Earth compared to KamLAND. The large target mass of SK can increase the significance rapidly when approaching the CCSN, resulting in a much higher event rate. But due to its relatively high background rate, SK is less sensitive to small signals. By noting that these two detectors are complementary in pre-SN neutrino search, a joint search combining measurements from these two detectors should improve the current detection sensitivity. We show in Section 6 that the combined alert benefits from the advantages of these two detectors. We expect the complementary properties of the two detectors to create a synergistic bond, enhancing the sensitivity to pre-SN neutrino signals.
@@ 6.1 Statistical approach for the combined search
The purpose of the combined pre-SN alert system is to provide a semi-realtime result of an online search of pre-SN neutrino signals. The chosen strategy is to perform a test of significance every five minutes upon the observed numbers of candidates and the expected numbers of background events in KamLAND and SK. Both experiments are considered to be Poisson counting experiments. The following likelihood function is constructed, which is a product of two Poissonian likelihoods,
where is the observed number of events and
The term , being the expected number of events, consists of the parameter for the number of signal and the expected number of background . The subscripts SK and KL denote Super-Kamiokande and KamLAND, respectively.
@@
For any pre-SN neutrino model, the neutrino fluxes in both detectors should be the same, and thus there are correlations between numbers of signals in KamLAND and SK. However, we note again that the test statistic merely reflects the level of agreement between data and the background-only hypothesis. Either or not the data is consistent with a pre-SN neutrino model is not tested. Therefore, and are treated as two independent parameters, and the degree of freedom is considered as two.
@@
Likewise, reactor neutrino background in the two detectors is correlated. This correlation does not affect the statistical test, because the expected number of background is estimated from the data taken in the background time window and normalized to the analysis time window, based on the assumption that the background rate is consistent with the background data taken before. However, such an assumption may not hold, since the reactor background can change significantly within a week if nuclear reactors are turned on or off. An unexpected increase of reactor flux, for example, when several nearby reactors are turned on, may cause an excess of events in the detectors. Although the background rate in the high reactor activity scenario is far below the alert criteria, a higher baseline can increase the risk of sending a false alarm. The reactor neutrino background is irreducible, as it consists of true events and its energy range overlaps that of the signal. Concerning this issue, we perform frequent background measurements in both detectors.
In the combined statistical test, the extremeness of an observation is quantified by calculating the ratio of maximum likelihoods with and without imposing background-only scenario ():
The background uncertainties are small compared to the Poissonian fluctuations of the observations, and thus are neglected to reduce computation time. The more the observation disagrees with the background-only hypothesis, the larger is. Based on likelihood ratio ordering, the -value of the combined search is determined by calculating the probability to find a signal that yields a no smaller than the observed one, . The corresponding significance is calculated by Equation 2.
@@
As mentioned in Section 5.3, the frequency to find a single combined search at least as extreme as the observation, i.e. the false alarm rate, is utilized to report the results of the pre-SN neutrino search. Figure 12 shows an example contour of the false alarm rate resulting from toy Monte-Carlo simulation. The x-axis and y-axis are numbers of candidates observed in KamLAND (in 24 hours) and SK (in 12 hours), respectively. The color of each box indicates the false alarm rate level of the corresponding numbers of observed events. If the observation drops in the yellow regions, which indicates , there is no hint of a pre-SN neutrino emission. If the observation is in the blue regions, where , it is considered as a significant excursion of observed event rate which could be caused by pre-SN .
@@ 6.2 Combined sensitivity
Following the above mentioned statistical approach, we perform a joint sensitivity study based on the estimations of , , and presented in Section 5. Figure 13 shows the time evolution of the combined sensitivity of pre-SN neutrinos from a Betelgeuse-like star of 15 M⊙, assuming medium reactor activity. For the Patton model and normal mass ordering, the warning time is extended to 12.4 hours prior to the CCSN. Similar to the discussions in Section 5, we assessed the sensitivities for the two pre-SN models, the three reactor activity cases, and the two neutrino mass orderings. Table 2 summarizes the warning time of the combined alert and the individual alerts. These results indicate that the combined alert presents an excellent performance, because the warning times are longer compared to either of the individual alerts. It is important to note that, even in the high reactor activity case, the warning time is at least 2.2 hours before the CCSN.
@@
These results demonstrate the complementarity of the KamLAND and the SK-Gd detectors. Taking advantages of the low background rate of KamLAND and the large target mass of SK, the combined alert presents improvement in the regard of extending the warning time as well as the distance coverage. In light of this, a combined pre-SN neutrino alert system was developed, and will be discussed in Section 7.
@@ 7 Combined online search for pre-SN neutrinos
The workflow of the system is introduced in the following. The system receives from both detectors the numbers of pre-SN neutrino candidates and the expected numbers of background. Processed by the DAQ systems, events in the two detectors are selected by their own selection processes following the descriptions in Section 3.1 and Section 4.1. The individual pre-SN alert software of KamLAND (SK) then counts the number of observed candidates () within a 24-hour (12-hour) time window. The expected number of background () is estimated using data from a background time window of 90 (30) days, as described in Section 5.1. Validity of data is also taken into consideration. Detector status is monitored and marked by a status code. When a detector undergoes calibration work, a test run or shutdown, the status of the detector is marked as “abnormal” and the data from this period will be invalidated. In addition to the above situations, there may also be network connection problems which can delay the data transfer. Therefore, the differences between the current time and the time when data are processed will also be checked. The individual KamLAND and SK pre-SN alert systems gather the above information, and exchange them between the servers of KamLAND and SK, as illustrated in Figure 15. The update frequency of the input is once every 5 minutes for each of the experiments. These inputs will be processed by the combined pre-SN alert software, yielding a result of the combined pre-SN search. The result will be exported to users, and the alert decision will be made based on the result.
@@
On the two servers, identically functioning software for the combined pre-SN alert system are installed. If either one of the software pipelines fails, the other one can still output the search result and issue alerts. Figure 16 illustrates the workflow of this software. The software runs on a precise repeating schedule once every 5 minutes. Validity of inputs is first checked by examining the status of each detector and the timestamps of the inputs. If a detector has an abnormal status, or if data from it is delayed for over 15 minutes, input from this detector is invalidated. In this case, instead of exporting the combined search result, the system will output result based only on the valid input. If all of the inputs are invalid, the result is not applicable. Only when both KamLAND and SK are in normal status and the data are up-to-date, the system exports the result of the combined search.
@@
In order to find out the correspondent false alarm rate, the software loads three pre-calculated false alarm rate tables, for KamLAND-only, SK-only, and their combination. Once the software finds a , an alarm will be sent the the Gamma-ray Coordinates Network (GCN) via an email-based circular. In addition, a text file containing the false alarm rate, along with a timestamp, and a code that denotes whether the result is for KamLAND-only, SK-only or the combined search, is always available on the KamLAND server and the SK server.
@@
The above processes, called main processes, are identical in both servers, as pictured in the blue box with solid border in Figure 16. An additional process as shown in the orange box with dashed border is uniquely installed on the SK server. In this process, the false alarm rate tables are updated automatically upon any changes in the expected numbers of background. This process, typically takes 40 minutes, is in parallel with the main processes, in order not to delay the output of the results. Therefore, the pre-calculated false alarm rate tables do not always correspond to the current background values. However, we note that a significant change in the background rates within an hour is unusual, because the background rates are obtained from measurements of a specific time window long enough to mitigate the effects of statistical fluctuations.
@@
This alert system, operational since May 2023 and accessible to the public, is designed to promptly notify astronomers and particle physicists to maintain operational readiness at their observatories, ensuring they do not miss any impending supernova events. Users can acquire identical results contained in the above mentioned text file from either of the servers, and are encouraged to check the consistency of the results from the two servers before putting into scientific use. Directional information of pre-SN neutrinos is not available from either SK or KamLAND. More information can be found at https://www.lowbg.org/presnalarm/, which is the official website of the combined pre-SN alert system.
@@ 8 Conclusion
The combined pre-supernova alert system performs a joint statistics test based on the data from the KamLAND and the Super-Kamiokande detectors. It has been operational and accessible to the public since May 2023. Integrating the complementary properties of the two detectors, the combined alert shows improved sensitivity to pre-supernova neutrinos. An optimistic warning time of 12.4 hours is obtained, for the Patton model and normal ordering, 1.5 hours longer than the Super-Kamiokande alert and 6.3 hours longer than the KamLAND alert, with the medium reactor activity assumption. At the same background level, its distance coverage for 15 M⊙ progenitors is 510 pc, which exceeds those of the individual alerts. While doubling the neutrino fluxes from nearby reactors increases significantly the backgrounds and affects the sensitivity of both detectors individually, the combined alert remain sensitive to pre-supernova neutrino emission with an expected warning time of no less than 2.2 hours for a Betelgeuse-like pre-supernova star, sufficiently long to cover the latency due to data processing. In addition, the combined alert system reduces the dead time for pre-supernova neutrino detection, promoting continuous monitoring even if one of the detectors is temporarily offline. All of these demonstrate the benefits of having a combined search for pre-supernova neutrinos.

### v3

@@ 5.6 Discussion
The results shown above suggest that the two detectors, KamLAND and SK, have different advantages in pre-SN neutrino detection. For KamLAND, the low background rate is an advantage in terms of resolving a small signal. As shown in Figure 6 (a), for a Betelgeuse-like pre-SN candidate, the expected number of signal events exceed the background count even tens of hours prior to the CCSN. As a result, KamLAND can provide a warning for nearby supernovae earlier than SK. On the other hand, the number of signal events in KamLAND is limited by the target mass, making it hard to reach stars at far distances. SK is sensitive to CCSN candidates further away from Earth compared to KamLAND. The large target mass of SK can increase the significance rapidly when approaching the CCSN, resulting in a much higher event rate. But due to its relatively high background rate, SK is less sensitive to small signals. By noting that these two detectors are complementary in pre-SN neutrino searches, a joint search combining measurements from these two detectors should improve the current detection sensitivity. We show in Section 6 that the combined alert benefits from the advantages of these two detectors. We expect the complementary properties of the two detectors to create a synergistic bond, enhancing the sensitivity to pre-SN neutrino signals.
@@ 6.1 Statistical approach for the combined search
The purpose of the combined pre-SN alert system is to provide a semi-realtime result of an online search of pre-SN neutrino signals. The chosen strategy is to perform a test of significance every five minutes upon the observed numbers of candidates and the expected numbers of background events in KamLAND and SK. The following likelihood function is constructed, which is a product of the Poisson likelihood of each experiment,
The test statistic based on likelihood ratio can be calculated by substituting for in Equation 3. The corresponding significance is found by calculating the chi-square quantile for two degrees of freedom.
@@
For any pre-SN neutrino model, the neutrino fluxes in both detectors should be the same, and thus there are correlations between numbers of signals in KamLAND and SK. However, we note that the test statistic merely reflects the level of agreement between data and the background-only hypothesis. Whether the data is consistent with a pre-SN neutrino model is not tested. Therefore, the parameters of signal in SK and KamLAND ( and ) are treated as two independent parameters, and the degrees of freedom is considered as two.
@@
Likewise, reactor neutrino background in the two detectors is correlated. This correlation does not affect the statistical test, because the expected number of background is estimated from the data taken in the background time window and normalized to the analysis time window, based on the assumption that the background rate is consistent with the background data taken before. However, such an assumption may not hold, since the reactor background can change significantly within a week if nuclear reactors are turned on or off. An unexpected increase of reactor flux, for example, when several nearby reactors are turned on, may cause an excess of events in the detectors. Although the background rate is still far below the alert criteria even in an extreme case we can imagine, i.e. the high reactor activity scenario, a higher baseline can increase the risk of sending a false alarm. The reactor neutrino background is irreducible, as it consists of true events and its energy range overlaps that of the signal. Concerning this issue, we perform frequent background measurements in both detectors.
@@
As mentioned in Section 5.3, the frequency to find a single combined search at least as extreme as the observation, i.e. the false alarm rate, is utilized to report the results of the pre-SN neutrino search. Figure 12 shows an example contour of the false alarm rate resulting from toy Monte-Carlo simulation. The x-axis and y-axis are numbers of candidates observed in KamLAND (in 24 hours) and SK (in 12 hours), respectively. The color of each box indicates the false alarm rate level of the corresponding numbers of observed events. If the observation drops in the yellow regions, which indicates a , there is no hint of a pre-SN neutrino emission. If the observation is in the blue regions, where the , it is considered as a significant excursion of the observed event rate which could be caused by pre-SN .
@@ 6.2 Combined sensitivity
Following the above mentioned statistical approach, we perform a joint sensitivity study based on the estimations of , , and presented in Section 5. Figure 13 shows the time evolution of the combined sensitivity of pre-SN neutrinos from a Betelgeuse-like star of 15 M⊙, assuming medium reactor activity. For the Patton model and normal mass ordering, the warning time is extended to 12.4 hours prior to the CCSN. Similar to the discussions in Section 5, we assessed the sensitivities for the two pre-SN models, the three reactor activity cases, and the two neutrino mass orderings. Table 2 summarizes the warning time of the combined alert and the individual alerts. These results indicate that the combined alert presents an improved performance, because the warning times are longer compared to either of the individual alerts. It is important to note that, even in the high reactor activity case, the warning time is at least 2.2 hours before the CCSN.
@@
These results demonstrate the complementarity of the KamLAND and the SK-Gd detectors. Taking advantage of the low background rate of KamLAND and the large target mass of SK, the combined alert presents improvement in extending the warning time as well as the distance coverage. In light of this, a combined pre-SN neutrino alert system was developed, and will be discussed in Section 7.
@@ 7 Combined online search for pre-SN neutrinos
The workflow of the system is introduced in the following. The system receives from both detectors the number of pre-SN neutrino candidates and the expected number of background. Processed by the DAQ systems, events in the two detectors are selected by their own selection processes following the descriptions in Section 3.1 and Section 4.1. The individual pre-SN alert software of KamLAND (SK) then counts the number of observed candidates () within a 24-hour (12-hour) time window. The expected number of background () is estimated using data from a background time window of 90 (30) days, as described in Section 5.1. Validity of data is also taken into consideration. Detector status is monitored and marked by a status code. When a detector undergoes calibration work, a test run or shutdown, the status of the detector is marked as “abnormal” and the data from this period will be invalidated. In addition to the above situations, there may also be network connection problems which can delay the data transfer. Therefore, the differences between the current time and the time when data are processed will also be checked. The individual KamLAND and SK pre-SN alert systems gather the above information, and exchange them between the servers of KamLAND and SK, as illustrated in Figure 15. The update frequency of the input is once every 5 minutes for each of the experiments. These inputs will be processed by the combined pre-SN alert software, yielding a result of the combined pre-SN search. The result will be exported to users, and the alert decision will be made based on the result.
@@
On the two servers, identically functioning software for the combined pre-SN alert system are installed. If either one of the software pipelines fails, the other one can still output the search result and issue alerts. Figure 16 illustrates the workflow of this software. The software runs on a precise repeating schedule once every 5 minutes. Validity of inputs is first checked by examining the status of each detector and the timestamps of the inputs. If a detector has an abnormal status, or if data from it is delayed for over 15 minutes, input from this detector is invalidated. In this case, instead of exporting the combined search result, the system will output a result based only on the valid input. If all of the inputs are invalid, the result is not applicable. Only when both KamLAND and SK are in normal status and the data are up-to-date, the system exports the result of the combined search.
@@
In order to determine the corresponding false alarm rate, the software loads three pre-calculated false alarm rate tables, for KamLAND-only, SK-only, and their combination. Once the software finds a , an alarm will be sent to the Gamma-ray Coordinates Network (GCN) via an email-based circular. In addition, a text file containing the false alarm rate, along with a timestamp, and a code that denotes whether the result is for KamLAND-only, SK-only or the combined search, is available to users who have registered on the official website of the combined pre-SN alert system (https://www.lowbg.org/presnalarm/).
@@
The above processes, called main processes, are identical in both servers, as pictured in the blue box with solid border in Figure 16. An additional process as shown in the orange box with dashed border is uniquely installed on the SK server. In this process, the false alarm rate tables are updated automatically upon any changes in the expected numbers of background. This process, typically takes 40 minutes, and is in parallel with the main processes, in order not to delay the output of the results. Therefore, the pre-calculated false alarm rate tables do not always correspond to the current background values. However, we note that a significant change in the background rates within an hour is unusual, because the background rates are obtained from measurements of a specific time window long enough to mitigate the effects of statistical fluctuations.
@@
This alert system, operational since May 2023 and accessible to the public, is designed to promptly notify astronomers and particle physicists to maintain operational readiness at their observatories, ensuring they do not miss any impending supernova events. Users can acquire identical results contained in the above mentioned text file from either of the servers, and are encouraged to check the consistency of the results from the two servers before putting into scientific use. Directional information of pre-SN neutrinos is not available from either SK or KamLAND. More information can be found on the official website (https://www.lowbg.org/presnalarm/).
@@ 8 Conclusion
The combined pre-supernova alert system performs a joint statistics test based on the data from the KamLAND and the Super-Kamiokande detectors. It has been operational and accessible to the public since May 2023. Integrating the complementary properties of the two detectors, the combined alert shows improved sensitivity to pre-supernova neutrinos. An optimistic warning time of 12.4 hours is obtained, for the Patton model and normal ordering, 1.5 hours longer than the Super-Kamiokande alert and 6.3 hours longer than the KamLAND alert, with the medium reactor activity assumption. At the same background level, its distance coverage for 15 M⊙ progenitors is 510 pc, which exceeds those of the individual alerts. While doubling the neutrino fluxes from nearby reactors increases significantly the backgrounds and affects the sensitivity of both detectors individually, the combined alert remains sensitive to pre-supernova neutrino emission with an expected warning time of no less than 2.2 hours for a Betelgeuse-like pre-supernova star, sufficiently long to cover the latency due to data processing. In addition, the combined alert system reduces the dead time for pre-supernova neutrino detection, promoting continuous monitoring even if one of the detectors is temporarily offline. All of these demonstrate the benefits of having a combined search for pre-supernova neutrinos.